# Why Machine Learning Models Systematically Underestimate Extreme Values

### v1

@@ 4 Discussion and Implications
This study examines the fundamental nature of attenuation bias in astronomical machine learning applications, demonstrating that when directly mapping from input features to output labels—a common practice in machine learning—this bias is inherent and widespread. The challenge stems from how modern machine learning methods, especially neural networks, typically handle uncertainties. While output uncertainties can be incorporated into loss (objective) functions, the treatment of input uncertainties remains challenging within most standard frameworks. This stands in contrast to traditional Bayesian approaches where input uncertainties can be explicitly considered in the analysis, although such treatment of input uncertainties is also often overlooked in practice in Bayesian inference applications to astronomical studies.
@@
Our analysis centers on linear regression, a choice motivated by two key considerations. First, linear regression permits analytical prediction of attenuation bias under certain assumptions, providing concrete mathematical insights into the effects of measurement uncertainties. Second, despite its simplicity, linear regression serves as a foundation for many astronomical analysis methods. Many power-law relations in astronomy (e.g., Kennicutt-Schmidt relation, Tully-Fisher relation, M-$\sigma$ relation) can be transformed into linear regressions in logarithmic space. Furthermore, many classical machine learning approaches are effectively generalized linear regression with extended, curated features, and neural networks can be viewed as a combination of a feature transformer and linear regression. Indeed, we demonstrated analytically that superlinear relationships typically suffer from even more severe attenuation bias.
@@ 4.1 Why Attenuation Bias Particularly Matters for Astronomy
While attenuation bias receives relatively little attention in general machine learning applications, it poses a particular challenge for astronomy due to the inherent dynamic range of astronomical measurements. For univariate or independent multivariate cases, the relevant scale is the ratio of signal variance to measurement error ($\sigma_{\text{range}}/\sigma_{x}$). The theoretical framework predicts that even at a ratio of 10—meaning the dynamic range of the data is approximately tenfold the measurement error—coefficient and label bias manifests at the percent level. This roughly corresponds to a SNR of 10. While general machine learning applications typically operate with much higher SNR (e.g., natural image processing), astronomy frequently operates in this critical regime where measurement uncertainties are significant relative to the signal range.
@@
Similar examples appear throughout astronomy. Stellar age measurements, except for the small subset of stars with precise asteroseismology, typically have uncertainties of $\mathcal{O}(1)$ Gyr—approximately 10% of the Hubble time (Lebreton & Goupil, 2014; Aerts, 2015; Valle et al., 2015; Xiang et al., 2017; Wu et al., 2018; Valentini et al., 2019; Claytor et al., 2020; Hayden et al., 2022). Elemental abundance measurements, even in dedicated studies, typically achieve precisions of 0.05-0.1 dex, again about 10% of the full dynamic range for [X/H] (Feltzing & Gustafsson, 1998; Carretta et al., 2002; Kirby et al., 2008; Xiang et al., 2017; Pancino et al., 2010; Holtzman et al., 2018).
The situation becomes particularly relevant for power-law relationships when transformed into log-log space for linear regression. For instance, supermassive black hole mass measurements spanning three-fiver orders of magnitude can only tolerate uncertainties of about 0.3-0.5 dex (e.g. Merritt & Ferrarese, 2001; Graham, 2012; Shen, 2013; Li et al., 2023)—corresponding to roughly a factor of three in linear scale—to avoid significant attenuation bias. Similar precision limitations appear in measurements of stellar mass and gas mass, where $\mathcal{O}(0.1)$ dex or more uncertainties are common in multi-dex power law studies (Conroy et al., 2009; Swindle et al., 2011; Leja et al., 2017, 2020). Thus, attenuation bias systematically affects astronomical measurements exactly where precision matters most.
@@ 4.2 Implications of Attenuation Bias for Spectral Analysis with Machine Learning
While this paper’s primary goal is to provide a general framework for understanding attenuation bias—which can be partially predicted in the simplistic limit of linear regression—we showcase a specific application: spectral analysis with machine learning. We chose this example for several compelling reasons. First, as noted above, the ratio between flux measurement error and flux variation naturally falls in the regime where attenuation bias becomes significant. Second, spectra provide an ideal test case for the theoretical framework, offering naturally correlated multivariate features that allow us to extend our analysis beyond univariate cases. Third, spectroscopy represents one of astronomy’s most important observational modes, yet remains largely unexplored in general machine learning literature, which has led the astronomical community to independently develop numerous specialized machine learning applications for spectral analysis (Fabbro et al., 2018; Leung & Bovy, 2019; Ting et al., 2019; Guiglion et al., 2020).
This interest has been further fueled by the need to analyze vast datasets of low-resolution, often low signal-to-noise ratio spectra from surveys such as LAMOST, DESI, and 4MOST. The prevalence of low-resolution data stems largely from modern spectroscopic surveys being cosmology-driven, where high resolution, while crucial for stellar studies, is less critical for galaxy observations.
@@
The challenge extends beyond traditional spectroscopy to even lower-resolution regimes, including narrow-band filters like the Gaia XP spectra and ground-based surveys such as J-PLUS (Cenarro et al., 2019; Wang et al., 2022a), S-PLUS (Mendes de Oliveira et al., 2019), and JPAS (Bonoli et al., 2021; Yuan et al., 2023). Given the persistent challenges in generating robust ab initio spectral models and the systematic uncertainties in theoretical models at these resolutions, researchers increasingly favor data-driven approaches, using high-resolution spectral observations or other precise measurements (from eclipsing binaries, astrometry, asteroseismology, or high-quality spectra) as training labels for machine learning models (Süveges et al., 2017; Ambrosch et al., 2023; Bello-García et al., 2023).
@@
While it is often argued that machine learning models can effectively transfer and match labels across different surveys, our study reveals potential problems with this assumption. Even for elemental abundances under optimal conditions—APOGEE resolution and SNR—attenuation bias remains non-negligible, compressing the dynamic range of inferred labels away from their original scale. Although this bias depends on the number of features employed and the range of stellar parameters (e.g., effective temperature as a nuisance parameter), and strongly correlated features can help minimize the impact, such mitigation is only possible for atomic species with numerous spectral features. Key elements, including widely studied species like oxygen, have limited atomic features in the optical wavelength ranges (Prakapavičius et al., 2013; Ting et al., 2018), leading to potentially severe attenuation (e.g. fig 13 in Xiang et al. 2019, fig 4 in Zhang et al. 2024). This effect is frequently observed in label transfer applications, particularly at low resolution, where one-to-one comparisons between training labels and predictions show more prominent skewing for elements with weaker and/or fewer features (e.g., figure 9-11 Leung & Bovy, 2019).
@@
This attenuation bias can have important implications, particularly for fundamental measurements of Galactic structure. For example, spectra contain information about stellar parameters that, when combined with apparent magnitudes, enable estimation of absolute magnitudes and hence distances—a variation of spectroscopic-photometric distance determination. The systematic underestimation of distances due to attenuation bias, which to our knowledge has not been well recognized or treated, can have cascading effects on our understanding of Galactic dynamics and chemical evolution.
@@
When these attenuated distances are used to map the Milky Way’s rotation curve, the radius $R$ at which the orbital velocity $v_{c}$ is measured is systematically underestimated. Given that the enclosed mass within radius $R$ is determined by $M(<R)=\frac{v_{c}^{2}R}{G}$, the percentage attenuation in distance measurements directly translates to the same percentage underestimation in the enclosed mass $M(<R)$. As we have established in this study, attenuation bias in low-resolution surveys can be significant even under conservative estimates, reaching the percent level. This systematic underestimation of the Milky Way’s mass has far-reaching implications for our understanding of the Galactic rotation curve, dark matter distribution, and overall Galactic dynamics. These findings suggest the need to revisit data-driven spectroscopic-photometric distance determinations in studies of the Milky Way.
@@ 4.3 Mitigation Strategies for Attenuation Bias
Given the significant implications of attenuation bias, it is natural to consider potential mitigation strategies. The most obvious approach involves calibration, either theoretical or empirical, acknowledging that this fundamental bias stems naturally from input uncertainty and is independent of training sample size and label accuracy. Importantly, the presence of this bias does not indicate that our machine learning models are incorrect or inadequately trained.
Yet, theoretical calibration through analytic calculations has its own limitations. While in this study we have shown that bias can be accurately predicted for univariate cases with homogeneous errors when uncertainties are well-measured, real data typically exhibits heteroskedastic (inhomogeneous) errors, making theoretical calculations substantially more complex. Although such calculations might be warranted for high-stakes studies of fundamental one-dimensional relationships, e.g., M-$\sigma$ (Colless et al., 1999; Gültekin et al., 2009; Sani et al., 2011; Zubovas & King, 2012; Shankar et al., 2016), Kennicutt-Schmidt relations (Robertson & Kravtsov, 2008; Schaye & Dalla Vecchia, 2008; Gnedin & Kravtsov, 2010; Onodera et al., 2010; Liu et al., 2011; Lada et al., 2013; Shi et al., 2018), their full development lies beyond the scope of this paper.
@@
Our study has shown that when features are not independent, their combined effect can help negate random input noise, thereby mitigating attenuation bias. This suggests that when possible, using data in its more raw form with multiple correlated estimators—each subject to independent noise—can be beneficial. For example, full spectral fitting based on individual pixels could help mitigate these issues compared to approaches using limited spectral indices or summary statistics that restrict the number of input features. Furthermore, more restricted models like linear regression might offer advantages in that their bias can be better predicted or calibrated, and they suffer less from attenuation compared to super-linear regression.
@@ 5 Conclusion
Acknowledgments
I am grateful to my colleagues at OSU whose lively discussions during our daily astro-coffee sessions have inspired some of this work. Special thanks to David Weinberg for his careful and invaluable feedback as this paper took shape. The proximity of our offices provided both intellectual stimulation and emotional urgency to complete this work. Some ideas in this research emerged from my ritual of bouncing thoughts off o1-preview and claude-3.5-sonnet - while their initial suggestions often led me down fascinating (and largely unproductive) rabbit holes, our back-and-forth exchanges proved invaluable. Alongside my human colleagues at OSU, I’d like to thank o1-preview and claude for their unwavering intellectual companionship during the course of this research.
@@ Appendix A Attenuation Bias Under Generalized Least Squares (GLS)
$y_{\text{true},i}=\beta x_{\text{true},i},$ (A1)
@@
$x_{\text{obs},i}$ $=x_{\text{true},i}+\delta_{x,i},$ (A2)
$y_{\text{obs},i}$ $=y_{\text{true},i}+\delta_{y,i},$ (A3)
@@
$\hat{\beta}_{GLS}=\frac{\sum_{i}w_{i}x_{\text{obs},i}y_{\text{obs},i}}{\sum_{i}w_{i}x_{\text{obs},i}^{2}}.$ (A4)
@@
$\hat{\beta}_{GLS}=\frac{\sum_{i}w_{i}(x_{\text{true},i}+\delta_{x,i})(\beta x_{\text{true},i}+\delta_{y,i})}{\sum_{i}w_{i}(x_{\text{true},i}+\delta_{x,i})^{2}}.$ (A5)
@@
$=\frac{\sum_{i}w_{i}\beta\sigma_{\text{range}}^{2}}{\sum_{i}w_{i}(\sigma_{\text{range}}^{2}+\sigma_{x}^{2})}.$ (A6)
@@
$\mathbb{E}[\hat{\beta}_{GLS}]=\beta\frac{\sigma_{\text{range}}^{2}}{\sigma_{\text{range}}^{2}+\sigma_{x}^{2}}.$ (A7)
@@
To demonstrate, we repeat the simulations described in the main text (see Section 2.1 and Figure 3), but now using GLS instead of OLS. The simulation setup remains largely the same, with a two-dimensional parameter space defined by the ratio of signal range to measurement error ($\sigma_{\text{range}}/\sigma_{x}$) and the measurement uncertainty in the dependent variable ($\sigma_{y}$). The key difference lies in the regression method: we now employ GLS, with weights set to $1/\sigma_{y}^{2}$ to account for the known variances in $y$.
@@
The GLS-based simulation results, shown in Figure 12, reinforce the conclusions drawn from the OLS case. Despite GLS accounting for known variances in $y$ through appropriate weighting, the results closely resemble those obtained using OLS. The attenuation factor $\lambda_{y}$ exhibits the same dependence on $\sigma_{\text{range}}/\sigma_{x}$, with no discernible influence from $\sigma_{y}$. This confirms that GLS does not mitigate the fundamental attenuation bias introduced by measurement errors in the independent variable $x$.
Figure 12: Similar to Figure 3, which presented ordinary least squares (OLS) results, this figure shows the two-dimensional attenuation factor $\lambda_{y}$ as a function of $\sigma_{\mathrm{range}}/\sigma_{x}$ (horizontal axis) and $\sigma_{y}$ (vertical axis) using Generalized Least Squares (GLS). Although GLS accounts for known variances in $y$ by adjusting the weights, the resulting attenuation factor maps remain effectively unchanged compared to OLS. This confirms that simply weighting the dependent variable according to its uncertainty does not mitigate the fundamental attenuation bias caused by measurement errors in the independent variable $x$. The color scale and contours are identical to those in Figure 3, enabling direct comparison.

### v2

@@ 4 Discussion and Implications
This study examines the fundamental nature of attenuation bias in astronomical machine learning applications, demonstrating that when directly mapping from input features to output labels—a common practice in machine learning—this bias is inherent and widespread. The challenge stems from how modern machine learning methods, especially neural networks, typically handle uncertainties. While output uncertainties can be incorporated into loss (objective) functions, the treatment of input uncertainties remains challenging within most standard frameworks. This stands in contrast to traditional Bayesian approaches where input uncertainties can be explicitly considered in the analysis.
@@
Our analysis centers on linear regression, a choice motivated by two key considerations. First, linear regression permits analytical prediction of attenuation bias under certain assumptions, providing concrete mathematical insights into the effects of measurement uncertainties. Second, despite its simplicity, linear regression serves as a foundation for many astronomical analysis methods. Many power-law relations in astronomy (e.g., Kennicutt-Schmidt relation, Tully-Fisher relation, M-$\sigma$ relation) can be transformed into linear regressions in logarithmic space. Furthermore, many classical machine learning approaches are effectively generalized linear regression with extended, curated features, and neural networks can be viewed as a combination of a feature transformer and linear regression.
@@ 4.1 Why Attenuation Bias Particularly Matters for Astronomy
While attenuation bias receives relatively little attention in general machine learning applications, it poses a particular challenge for astronomy due to the inherent dynamic range of astronomical measurements. For univariate or independent multivariate cases, the relevant scale is the ratio of signal variance to measurement error ($\sigma_{\text{range}}/\sigma_{x}$). The theoretical framework predicts that even at a ratio of 10—meaning the dynamic range of the data is approximately tenfold the measurement error—coefficient and label bias manifests at the percent level. While general machine learning applications typically operate with much higher SNR (e.g., natural image processing), astronomy frequently operates in this critical regime where measurement uncertainties are significant relative to the signal range.
@@
Similar examples appear throughout astronomy. Stellar age measurements, typically have uncertainties of $\mathcal{O}(1)$ Gyr—approximately 10% of the Hubble time (Lebreton and Goupil, 2014; Aerts, 2015; Valle et al., 2015; Xiang et al., 2017; Wu et al., 2018; Valentini et al., 2019; Claytor et al., 2020; Hayden et al., 2022). Elemental abundance measurements, typically achieve precisions of 0.05-0.1 dex, again about 10% of the full dynamic range for [X/H] (Feltzing and Gustafsson, 1998; Carretta et al., 2002; Kirby et al., 2008; Xiang et al., 2017; Pancino et al., 2010; Holtzman et al., 2018).
The situation becomes particularly relevant for power-law relationships when transformed into log-log space for linear regression. For instance, supermassive black hole mass measurements spanning three-fiver orders of magnitude can only tolerate uncertainties of about 0.3-0.5 dex (e.g. Merritt and Ferrarese, 2001; Graham, 2012; Shen, 2013; Li et al., 2023)—corresponding to roughly a factor of three in linear scale—to avoid significant attenuation bias. Similar precision limitations appear in measurements of stellar mass and gas mass, where $\mathcal{O}(0.1)$ dex or more uncertainties are common in multi-dex power law studies (Conroy et al., 2009; Swindle et al., 2011; Leja et al., 2017, 2020). Thus, attenuation bias systematically affects astronomical measurements exactly where precision matters most.
@@ 4.2 Implications of Attenuation Bias for Spectral Analysis with Machine Learning
While this paper’s primary goal is to provide a general framework for understanding attenuation bias—which can be partially predicted in the simplistic limit of linear regression—we showcase a specific application: spectral analysis with machine learning. We chose this example for several compelling reasons. First, as noted above, the ratio between flux measurement error and flux variation naturally falls in the regime where attenuation bias becomes significant. Second, spectra provide an ideal test case for the theoretical framework, offering naturally correlated multivariate features that allow us to extend our analysis beyond univariate cases. Third, spectroscopy represents one of astronomy’s most important observational modes, yet remains largely unexplored in general machine learning literature, which has led the astronomical community to independently develop numerous specialized machine learning applications for spectral analysis (Fabbro et al., 2018; Leung and Bovy, 2019; Ting et al., 2019; Guiglion et al., 2020).
@@
The challenge extends beyond traditional spectroscopy to even lower-resolution regimes, including narrow-band filters like the Gaia XP spectra and ground-based surveys such as J-PLUS (Cenarro et al., 2019; Wang et al., 2022b), S-PLUS (Mendes de Oliveira et al., 2019), and JPAS (Bonoli et al., 2021; Yuan et al., 2023). Given the persistent challenges in generating robust ab initio spectral models, researchers increasingly favor data-driven approaches, using high-resolution spectral observations or other precise measurements (from eclipsing binaries, astrometry, asteroseismology, or high-quality spectra) as training labels for machine learning models (Süveges et al., 2017; Ambrosch et al., 2023; Bello-García et al., 2023).
@@
While it is often argued that machine learning models can effectively transfer and match labels across different surveys, our study reveals potential problems with this assumption. Even for elemental abundances under optimal conditions—APOGEE resolution and SNR—attenuation bias remains non-negligible, compressing the dynamic range of inferred labels away from their original scale. Although this bias depends on the number of features employed and the range of stellar parameters (e.g., effective temperature as a nuisance parameter), and strongly correlated features can help minimize the impact, such mitigation is only possible for atomic species with numerous spectral features. Key elements, including widely studied species like oxygen, have limited atomic features in the optical wavelength ranges (Prakapavičius et al., 2013; Ting et al., 2018), leading to potentially severe attenuation (e.g. fig 13 in Xiang et al. 2019, fig 4 in Zhang et al. 2024). This effect is frequently observed in label transfer applications, particularly at low resolution, where one-to-one comparisons between training labels and predictions show more prominent skewing for elements with weaker and/or fewer features (e.g., figure 9-11 Leung and Bovy, 2019).
@@
This attenuation bias can have important implications, particularly for fundamental measurements of Galactic structure. For example, spectra contain information about stellar parameters that, when combined with apparent magnitudes, enable estimation of absolute magnitudes and hence distances—a variation of spectroscopic-photometric distance determination. The systematic underestimation of distances due to attenuation bias, whic to our knowledge has not been well recognized or treated, can have cascading effects on our understanding of Galactic dynamics and chemical evolution.
@@
When these attenuated distances are used to map the Milky Way’s rotation curve, the radius $R$ at which the orbital velocity $v_{c}$ is measured is systematically underestimated. Given that the enclosed mass within radius $R$ is determined by $M(<R)=v_{c}^{2}R/G$, the percentage attenuation in distance measurements directly translates to the same percentage underestimation in the enclosed mass $M(<R)$. As we have established in this study, attenuation bias in low-resolution surveys can be significant even under conservative estimates, reaching the percent level. This systematic underestimation of the Milky Way’s mass has far-reaching implications for our understanding of the Galactic rotation curve, dark matter distribution, and overall Galactic dynamics. These findings suggest the need to revisit data-driven spectroscopic-photometric distance determinations in studies of the Milky Way.
@@ 4.3 Mitigation Strategies for Attenuation Bias
Given the significant implications of attenuation bias, it is natural to consider potential mitigation strategies. The most obvious approach involves calibration, either theoretical or empirical, acknowledging that this fundamental bias stems naturally from input uncertainty and is independent of training sample size and label accuracy.
Yet, theoretical calibration through analytic calculations has its own limitations. While in this study we have shown that bias can be accurately predicted for univariate cases with homogeneous errors when uncertainties are well-measured, real data typically exhibits heteroskedastic (inhomogeneous) errors, making theoretical calculations substantially more complex.
@@
Our study has shown that when features are not independent, their combined effect can help negate random input noise, thereby mitigating attenuation bias. This suggests that when possible, using data in its more raw form with multiple correlated estimators—each subject to independent noise—can be beneficial. For example, full spectral fitting based on individual pixels could help mitigate these issues compared to approaches using limited spectral indices or summary statistics that restrict the number of input features. Furthermore, more restricted models like linear regression might offer advantages in that their bias can be better predicted or calibrated, and they suffer less from attenuation compared to super-linear regression. As demonstrated in this study, linear regression permits analytical prediction of attenuation bias under certain assumptions, whereas for super-linear regression, such theoretical predictions become intractable.
@@ 5 Conclusion
I am grateful to my colleagues at OSU whose lively discussions during our daily astro-coffee sessions have inspired some of this work. Special thanks to David Weinberg for his careful and invaluable feedback as this paper took shape. The proximity of our offices provided both intellectual stimulation and emotional urgency to complete this work. Some ideas in this research emerged from my ritual of bouncing thoughts off o1-preview and claude-3.5-sonnet - while their initial suggestions often led me down fascinating (and largely unproductive) rabbit holes, our back-and-forth exchanges proved invaluable. Alongside my human colleagues at OSU, I’d like to thank o1-preview and claude for their unwavering intellectual companionship during the course of this research. This research is supported by the National Science Foundation under Grant No. AST-2406729.
@@ Appendix A Attenuation Bias Under Generalized Least Squares (GLS)
$y_{\text{true},i}=\beta x_{\text{true},i},$ (46)
@@
$x_{\text{obs},i}$ $=x_{\text{true},i}+\delta_{x,i},$ (47)
$y_{\text{obs},i}$ $=y_{\text{true},i}+\delta_{y,i},$ (48)
@@
$\hat{\beta}_{GLS}=\frac{\sum_{i}w_{i}x_{\text{obs},i}y_{\text{obs},i}}{\sum_{i}w_{i}x_{\text{obs},i}^{2}}.$ (49)
@@
$\hat{\beta}_{GLS}=\frac{\sum_{i}w_{i}(x_{\text{true},i}+\delta_{x,i})(\beta x_{\text{true},i}+\delta_{y,i})}{\sum_{i}w_{i}(x_{\text{true},i}+\delta_{x,i})^{2}}.$ (50)
@@
$=\frac{\sum_{i}w_{i}\beta\sigma_{\text{range}}^{2}}{\sum_{i}w_{i}(\sigma_{\text{range}}^{2}+\sigma_{x}^{2})}.$ (51)
@@
$\mathbb{E}[\hat{\beta}_{GLS}]=\beta\frac{\sigma_{\text{range}}^{2}}{\sigma_{\text{range}}^{2}+\sigma_{x}^{2}}.$ (52)
@@
To demonstrate, we repeat the simulations described in the main text (see Section 2.1), but now using GLS instead of OLS. The simulation setup remains largely the same, with a two-dimensional parameter space defined by the ratio of signal range to measurement error ($\sigma_{\text{range}}/\sigma_{x}$) and the measurement uncertainty in the dependent variable ($\sigma_{y}$). The key difference lies in the regression method: we now employ GLS, with weights set to $1/\sigma_{y}^{2}$ to account for the known variances in $y$.
@@
Our numerical simulations further confirm that the attenuation factor $\lambda_{y}$ depends only on $\sigma_{\text{range}}/\sigma_{x}$ and is independent of $\sigma_{y}$, even when using GLS regression. This demonstrates that GLS does not mitigate the fundamental attenuation bias introduced by measurement errors in the independent variable $x$.